OpenStudy (anonymous):

rotation 90 degrees counterclockwise about the origin

OpenStudy (rsadhvika):

\[\large \left(\begin{matrix}x'\\y'\end{matrix}\right) = \left[\begin{matrix}0&-1\\1&0\end{matrix}\right]\left(\begin{matrix}x\\y\end{matrix}\right)\]

OpenStudy (rsadhvika):

\(\large (x, ~y)\leadsto (-y, ~x)\)